the ratio between two complementary is 7 to 2. find the measure of the smaller angle
Answer by stanbon(75887)

the ratio between two complementary is 7 to 2. find the measure of the smaller angle
----
1st angle: x
2nd angle: 90-x
----
Equation:
x/(90-x) = 7/2
2x = 630-7x
9x = 630
x = 70 (1st angle)
90-x = 20 (2nd angle)
=========================
Cheers,
Stan H.